### v2

@@ 4.1.2 Local Error Analysis
Figure 12: Instantaneous mesh-based flow-fields on time-extrapolated snapshot (t=10.5s) for (a) Re=1600 and (b) Re=3200. From left-to-right: P=7 target solution (DNS), P=1 input coarse solution (projected DNS), spectral element interpolation, GNN in Model 1 configuration (coarse-scale) using 26 neighbors, GNN in Model 2 configuration (multiscale) using 26 neighbors. Top row shows z-component vorticity contours in 3D (blue and red indicate vorticity of 3 and -3 respectively), middle row shows z-component of vorticity in the X-Y plane at z=$\pi/2$, and bottom row shows velocity magnitude in the same 2D plane. Black and white circles in 2D visualizations are provided as visualization guidelines.
4.2 Reynolds Number Extrapolation
@@
To build on the above analysis, which focused on pure time-extrapolation of models, the objective of this section is to investigate model performance in Reynolds extrapolated regimes. In other words, global and local error quantities are investigated with the goal of assessing how a particular GNN model trained using Re=1600 data, for example, performs when predicting super-resolved fields at Re=3200 (and vice-versa). At the Reynolds numbers considered here for the TGV configuration, combined with the inherent challenge of the one-shot P=1-to-7 task, these extrapolations are difficult by design for any data-based model – as such, comparisons between Reynolds-extrapolated and non-extrapolated predictions facilitates a better understanding of model capabilities and generalization limits, which in turn leads to useful physical interpretations of the mechanisms by which these GNN models generate their predictions.
@@
A global perspective of Reynolds-extrapolated predictions is provided in the energy spectrum plots in Fig. 13, with Fig. 13(a) showing results for Re=1600 and Fig. 13(b) for Re=3200. More specifically, Fig. 13(a) compares super-resolution predictions using an input P=1 field at Re=1600 for two GNNs: one trained at the same Re=1600, and the other trained using Re=3200 data. The analog is true for Fig. 13(b). As a result, Fig. 13(a) illustrates the impact of downward/reduced Reynolds extrapolation by a factor of 2 on the predicted spectrum, and Fig. 13(b) illustrates the same, but for an upward/increased Reynolds number extrapolation. It should be noted that for ease of visualization, comparisons between GNNs in both Model 1 and 2 configurations are shown at a coarse element neighborhood size of 26 – overall Reynolds number extrapolation tendencies of models were found to hold for the other neighborhood sizes.
@@
In the Re=1600 predictions shown in Fig. 13(a), the GNN extrapolations (corresponding to models trained at Re=3200) inject additional energy at the higher wavenumbers relative to the models trained at the same Reynolds number of 1600, resulting in higher levels of overshoot and spectrum relative errors in this regime. The notable increase in errors at the higher wavenumbers are balanced by decreases to errors in the medium-range wavenumbers. In the reverse extrapolation direction in Fig. 13(b), the upwards-extrapolated models (now corresponding to those trained at Re=1600) result in essentially the inverse trend as the extrapolations in Fig. 13(a): at Re=3200, both Models 1 and 2, relative to the non-extrapolated model, tend to produce lower energies relative to the target at larger wavenumbers. In Fig. 13(b), this results in a greater level of error at the lower P=1 Nyquist limit (k=36) than the non-extrapolated model, in direct contrast to the increase in error near the high Nyquist limit observed in Fig. 13(a). The spectra trends in the extrapolated regimes point to a quality of the super-resolution task that is inherent to this complex one-to-many mapping problem: since models were trained only at a single Reynolds number, their outputs for a given coarse-scale input are naturally ”hard-coded” to the particular Reynolds number at which they are trained on, resulting in the deviations observed in Fig. 13 (i.e., it is reasonable to expect a model purely trained at Re=3200 to overshoot energy when extrapolating to Re=1600, and a model purely trained at Re=1600 is expected to undershoot energy when extrapolated to Re=3200). In light of this implication and the deviations observed in Fig. 13, a more localized analysis of flow patterns in physical space is warranted.
@@
To this end, Fig. 14 shows the physical space analog of Fig. 13, providing corresponding visualizations of instantaneous super-resolved flow-fields. Similar to the previously discussed Fig. 12, in Fig. 14, target fields are compared with Model 2 outputs in non-extrapolated (third columns) and Reynolds-extrapolated (fourth columns, highlighted in red text) prediction settings on a test set snapshot. In both Re=1600 and Re=3200 cases (Fig. 14(a) and (b), respectively), the vorticity contour visualizations in the top rows reveal the ability of the GNN to recover much of the fine-scale turbulent structures even in extrapolated regimes. However, in the Re=1600 case (Fig. 14(a)), since the extrapolated predictions use the model trained on Re=3200 data, vorticity contours appear more ”dense” and noisy, indicative of the energy overshoot observed in Fig. 13(a) – the opposite is true for the Re=3200 case in Fig. 14(b), with extrapolated predictions sourced from Re=1600 models resulting in a less dense/rich (i.e., less turbulence intensity) super-resolved field relative to the target. The 2D planes of vorticity and velocity magnitude in the middle and bottom rows of Fig. 13, respectively, show how the downwards-extrapolation in Reynolds number (Fig. 13(a)) is in good qualitative agreement to the non-extrapolated model, and in-turn, the target. The same is not entirely true in the assessment of upwards-extrapolation in Reynolds number (Fig. 13(b)); although fine-scale structures are indeed generated by the Re=1600-trained model here, there is pronounced difficulty in producing all of the fine-scale structures as observed in the non-extrapolated model.
@@ 5 Conclusion
The goal of this work is to enable mesh-based super-resolution of fluid flows in 3D spatial configurations. A new graph neural network architecture is introduced to this end, which is designed to take as input a graph representation of a highly coarsened unresolved flow-field, and produce as output the corresponding super-resolved flow-field on a high-resolution graph in which fine-scale features exist. The graph generation procedure coincides with an element-local interpretation of flow-fields; it is consistent with strategies used in spectral element discretizations, and shares similarities with those used in finite element and finite volume discretizations. In this framework, the GNN is designed to operate not on the full mesh-based field at once, but on localized meshes of elements (or cells) directly. To facilitate super-resolution over element-local graphs, the architecture is composed of two key stages. In the first stage, a query element alongside a set number of neighboring coarse elements are encoded into a single latent graph using coarse-scale synchronized message passing over the element neighborhood (termed the coarse-scale processor). In the second stage, fine-scale message passing is carried out on the latent graph, and is used to correct for interpolation errors produced by a graph unpooling operation (termed the fine-scale processor).
@@
As a demonstration study, the architecture is applied in this work to achieve super-resolution in the Taylor-Green Vortex (TGV) configuration at Re=1600 and Re=3200, with hexahedral mesh-based data generated using the NekRS flow solver. Within this application, the goal is to move from a P=1 element discretization (highly coarsened projected DNS) to a P=7 (DNS-level) discretization.
@@
Through the multiscale GNN approach, specification of the number of message passing layers in either the coarse or fine-scale processors directly controls the length scales at which information exchange occurs throughout the super-resolution process, leading to valuable physical insights. As such, demonstrations of TGV super-resolution focused on performance assessments from two angles: (1) comparison of a purely coarse-scale model configuration (Model 1, with 0 message passing layers in the FSP) with a multiscale configuration (Model 2, with layers added to the FSP), and (2) the effect of the input coarse element neighborhood size (0, 6, and 26 neighbors) on the super-resolution accuracy. Although all configurations were found to competitively achieve super-resolution in time-extrapolated scenarios (with significant improvements obtained over spectral element interpolations), model errors were found to scale roughly in proportion to increase in Reynolds number, and accuracy improvements provided by changes to the model type and number of coarse element neighbors were also found to depend on the Reynolds number. For example, at Re=1600, jumping from 0 to 6 (and 26) neighbors improved results, with Model 2 achieving generally superior reconstruction quality over Model 1 counterparts. At Re=3200, performance between Models 1 and 2 were comparable, with minimal sensitivity to the number of neighbors (i.e., at Re=3200, information contained in the surrounding element neighborhood did not play much of a role in the super-resolution process). This finding led to a surprising conclusion: although Model 2 appears to be generally preferred over a wider range of Reynolds numbers, a competitive level of fine-scale information content can be produced by an element-local model reliant on learning purely coarse-scale message passing operations (the Model 1 configuration).
@@
The work concluded with demonstrations of performance at extrapolated Reynolds numbers. Two Reynolds-extrapolation scenarios were tested: downwards-extrapolation (application of a model trained at Re=3200 on Re=1600 data) and upwards-extrapolation (application of a model trained at Re=1600 on Re=3200 data). Although models retained their ability to generate fine-scale information at an appreciable level, upwards-extrapolation presented a greater challenge for the GNN models, with models performing better in downwards-extrapolation settings (although downwards Reynolds extrapolation resulted in overprediction of energy at the fine scales). This is consistent with physical expectations, and was deemed to be a consequence of the training procedure: the element-local flow-fields at Re=3200 naturally span a wider range of encountered turbulent structures relative to those observed at Re=1600, resulting in a highly challenging upwards-extrapolation task in the Reynolds number.
@@
This study presents many avenues for future work that can take direct advantage of the GNN architecture introduced here. As mentioned in Sec. 4.2, a promising extension to improve Reynolds number extrapolations is to source the training data from many different Reynolds numbers. Another pathway is to extend an incremental training procedure (explored to good effect in Ref. [62]) to the GNN setting, which decomposes the one-shot approach into a sequence of smaller, but easier, super-resolution increments. Due to the node permutation invariance properties provided by the fully graph-based architecture, the advantage in such an approach is that the same GNN model can be applied at many different resolution increments, potentially providing a more robust model. Additionally, within the one-shot context used here, investigation of the effect of model hyperparameters on generalization capability (such as the hidden channel dimensionality and number of message passing layers) can be explored in greater detail. Lastly, a particularly promising avenue is extending this architecture to flow configurations in more complex geometries exhibiting boundary layers and flow separation.

### v3

@@ 4.1.3 Physical Space Visualizations
4.2 Demonstration on Backward-Facing Step (BFS)
In this section, to augment the TGV analysis, super-resolution GNNs following the Model 2 (multiscale) setting trained independently on the BFS data are evaluated. For model assessment, this section considers BFS-trained GNN evaluations on unseen BFS test set snapshots at the same training Reynolds numbers (i.e., temporal extrapolation).
4.2.1 Global Error Analysis
Due to the non-cubic nature of the BFS domain, executing a spectrum based global error analysis in the Fourier basis is non-ideal. As a workaround, global errors are assessed using proper orthogonal decomposition (POD), which allows for the construction an orthogonal basis tailored to a specific geometry from a set of velocity field snapshots. More specifically, the POD basis vectors are extracted from target P=7 ground-truth flow-fields using the method of snapshots [79], which produces the basis using an eigendecomposition on the temporal correlation matrix. As this is a well-known approach, the finer details are not covered here – the interested reader is referred to C and Ref. [79] for method details.
Instantaneous BFS POD spectra for a testing set snapshot are shown in Fig. 13(a) and (b) for Re=1600 and 3200, respectively. Also included in the figure is a subset of the 20 total POD modes for the Re=1600 decomposition. Apparent at both Reynolds numbers is the lack of energy content in the P=1 coarse input – the downward shift of the P=1 energy content at all modes (blue curves in Fig. 3) reflects the expected elimination of all small-scale content in the flow. The relative error plots, analogous to the TGV spectrum relative errors shown earlier, highlight the GNN’s capability to recover a significant amount of fine-scale information content from coarse-scale inputs. More specifically, with the exception of mode index 0 at Re=1600, GNN predictions improve on the SE interpolation baseline across the board. The benefits incurred by moving from 0 to 26 coarse element neighbors in the GNN input is especially apparent in the relative error trends. Additionally, mirroring the TGV trends, relative errors in the GNN spectra increase when moving to Re=3200, and the reconstruction advantage provided by the 26-neighbor model appears to be diminishing with increasing Reynolds number.
Overall, the POD modes are quite difficult to interpret. However, visualization of the modes provide some avenue of interpretation for the global error reductions (or increases) provided by the GNNs at specific mode indices. Interestingly, mode index 0 is the only case at Re=1600 where the GNNs achieve higher relative error than the baseline SE interpolation. The POD mode at this index points to the activation of several large-scale features not present in the other modes, such as the top-wall boundary layer and the shear layer emerging from the separation point. The increase in error here could be indicative of the same relative error increase at the larger length scales (smaller wavenumbers) observed in the TGV spectra (Sec. 4.1). Despite this, errors at all other mode indices are quite substantially reduced relative to the SE interpolation – when observing some of the modes, it appears that these error reductions come from a more accurate reconstruction of (a) complex flow features near the reattachment point, and (b) downstream turbulence. At Re=1600, this is especially apparent for Mode 5, which shows a dramatic reduction of error when increasing the GNN coarse element neighbors from 0 to 26. This mode (alongside modes 2 and 18) contains small-scale effects near the step cavity, and highly complex small-scale features at the separation point and further downstream.
Figure 13: (a) POD energy spectrum (top plot) and relative errors in spectrum (bottom plot) evaluated at an unseen Re=1600 BFS snapshot. Curves shown for for target (black), coarse input (blue), SE interpolation (gray), GNN with 0 neighbors (dashed green), and GNN with 26 neighbors (dashed red). GNN correspond to Model 2 (multiscale) configuration. Snapshots at the bottom show visualizations of a subset of POD modes, indicated by black arrows. (b) Same as (a), but for Re=3200.
4.2.2 Physical Space Visualizations
Physical space reconstructions for the BFS models at both Re=1600 and 3200 are provided in Figs. 14 and 15, which show 3d visualizations of vorticity field and additional visualizations of velocity fields on extracted 2d X-Y and Y-Z planes.
Figure 14: Physical space visualizations of target (P=7), input (P=1), SE interpolation, and predicted GNN (Model 2, 26 neighbors) flow-fields for the BFS at unseen Re=1600 and 3200 snapshots. (a) Z-component vorticity contours in 3D (blue and red indicate vorticity of 5 and -5 respectively). (b) Z-component vorticity fields in the XY plane (taken at z=1, the spanwise midpoint). (c) Same as (b), but showing velocity magnitude fields.
Figure 15: Physical space visualizations of target (P=7), input (P=1), SE interpolation, and predicted GNN (Model 2, 26 neighbors) flow-fields for the BFS at unseen Re=1600 and 3200 snapshots (same as those used in Fig. 14) (a) Z-component vorticity fields in the YZ plane (taken near the streamwise location of reattachment point). (b) Same as (a), but showing velocity magnitude fields.
Inspection of the 3d vorticity contours in Fig. 14(a) shows similar qualitative behavior to the TGV trends. The GNNs (in this case, only Model 2 configurations are shown) add noticeable turbulent content to the flow from the P=1 input, resulting in expected improvement over the SE interpolation. Although the GNNs have difficulty capturing the correct vorticity contours upstream of the step anchor, the downstream turbulence enrichment is apparent, reflecting the spectrum error improvements described above. Although not shown in Fig. 14, improvement gains when jumping from 0 to 26 neighbors in the BFS GNNs come from better predictions in the upstream boundary layer, in the step cavity, and in the elimination of many nonphysical blocky artifacts in the downstream turbulence.
A more detailed inspection of the super-resolution predictions are provided in the form of 2d planes: an X-Y plane extracted at the spanwise mid-section of the domain (shown in Fig. 14), and a Y-Z plane extracted near the streamwise location of reattachment point (shown in Fig. 15).
Assessment of the XY planes reveal the primary streamwise-dominated flow features, such as the shear layer, separation point, and reattachment of the flow. At Re=1600, the GNN recovers a considerable amount of the vorticity content, and most crucially recovers a similar shear layer and upstream boundary layer thickness relative to the target, ”unsmearing“ these quantities from the coarse P=1 input, an effect that is not captured by the SE interpolation. At Re=3200, results mirror those encountered in the TGV, suggesting a much more difficult P=1 to 7 mapping task for the GNN. Despite the fact that voriticity intensities are not captured as well as the Re=1600 case, the GNN is able to super-resolve the primary flow features fairly well in the more challenging Re=3200 scenario, eliminating many of the blocky artifacts found in the SE interpolation. A similar story is seen in the YZ planes near the reattachment point, shown in Fig. 15, which highlights recovery of some of the span-wise (3D) effects – these planes in particular showcase the recovery of vorticity intensity and fine-scale velocity flow features in the Re=1600 case near the lower wall.
4.3 Assessment of Reynolds Number Extrapolation
@@
To build on the above analysis in Sec. 4.1 and 4.2, which focused on purely time-extrapolation of models, the objective of this section is to investigate model performance in Reynolds extrapolated regimes. In other words, global and local error quantities are investigated with the goal of assessing how a particular GNN model trained using Re=1600 data, for example, performs when predicting super-resolved fields at Re=3200 (and vice-versa). Reynolds extrapolation is shown only for the TGV case – general trends were found to be applicable across different configurations.
@@
A global perspective of Reynolds-extrapolated predictions is provided in the energy spectrum plots in Fig. 16. Figure 16(a) compares super-resolution predictions using an input P=1 field at Re=1600 for two GNNs: one trained at the same Re=1600, and the other trained using Re=3200 data. The analog is true for Fig. 16(b). As a result, Fig. 16(a) illustrates the impact of downward/reduced Reynolds extrapolation by a factor of 2 on the predicted spectrum, and Fig. 16(b) illustrates the same, but for an upward/increased Reynolds number extrapolation.
@@
In the Re=1600 predictions shown in Fig. 16(a), the GNN extrapolations (corresponding to models trained at Re=3200) inject additional energy at the higher wavenumbers relative to the models trained at the same Reynolds number of 1600, resulting in higher levels of overshoot and spectrum relative errors in this regime. The notable increase in errors at the higher wavenumbers are balanced by decreases to errors in the medium-range wavenumbers. In the reverse extrapolation direction in Fig. 16(b), the upwards-extrapolated models (now corresponding to those trained at Re=1600) result in essentially the inverse trend as the extrapolations in Fig. 16(a): at Re=3200, both Models 1 and 2, relative to the non-extrapolated model, tend to produce lower energies relative to the target at larger wavenumbers. In Fig. 16(b), this results in a greater level of error at the lower P=1 Nyquist limit (k=36) than the non-extrapolated model, in direct contrast to the increase in error near the high Nyquist limit observed in Fig. 16(a). The spectra trends in the extrapolated regimes point to a quality of the super-resolution task that is inherent to this complex one-to-many mapping problem: since models were trained only at a single Reynolds number, their outputs for a given coarse-scale input are naturally ”hard-coded” to the particular Reynolds number at which they are trained on, resulting in the deviations observed in Fig. 16 (i.e., it is reasonable to expect a model purely trained at Re=3200 to overshoot energy when extrapolating to Re=1600, and a model purely trained at Re=1600 is expected to undershoot energy when extrapolated to Re=3200).
In light of this implication and the deviations observed in Fig. 16, a more localized analysis of flow patterns in physical space is warranted.
@@
To this end, Fig. 17 shows the physical space analog of Fig. 16, providing corresponding visualizations of instantaneous super-resolved flow-fields. In both Re=1600 and Re=3200 cases (Fig. 17(a) and (b), respectively), the vorticity contour visualizations in the top rows reveal the ability of the GNN to recover much of the fine-scale turbulent structures even in extrapolated regimes. However, in the Re=1600 case (Fig. 17(a)), since the extrapolated predictions use the model trained on Re=3200 data, vorticity contours appear more ”dense” and noisy, indicative of the energy overshoot observed in Fig. 16(a) – the opposite is true for the Re=3200 case in Fig. 17(b), with extrapolated predictions sourced from Re=1600 models resulting in a less dense/rich (i.e., less turbulence intensity) super-resolved field relative to the target. The 2D planes of vorticity and velocity magnitude show how the downwards-extrapolation in Reynolds number (Fig. 16(a)) is in good qualitative agreement to the non-extrapolated model, and in-turn, the target. The same is not entirely true in the assessment of upwards-extrapolation in Reynolds number (Fig. 16(b)); although fine-scale structures are indeed generated by the Re=1600-trained model here, there is pronounced difficulty in producing all of the fine-scale structures as observed in the non-extrapolated model.
@@
4.4 Assessment of Geometry Extrapolation on the Cavity
In this section, the cavity configuration is used to assess the zero-shot geometry extrapolation capability of models. More specifically, in the text below, GNNs trained on both TGV and BFS configurations are used to super-resolve flow snapshots in an entirely different geometric configuration, namely the cavity flow. These evaluations are then compared to a model trained separately on cavity flow snapshots, which provides a performance upper bound on the TGV- and BFS-trained GNNs. To isolate purely geometry extrapolation behavior and errors, geometry extrapolation analysis is restricted to Re=1600.
Instantaneous cavity fields at Re=1600 are shown in Fig. 18, which provides XY planes overlaid with 3D vorticity contours (a view into the plane) of target, input, SE interpolation, and GNN-predicted flow-fields. The GNN predictions shown in Fig. 18 correspond to the Model 2 configuration. Similar to the BFS vorticity contour plots, the visualizations showcase the ability of the GNNs to generate much of the turbulence structure in the flow. This is apparent when comparing to the SE interp outputs, which showcase highly blocky and coarse artifacts, upon which the GNNs improve.
Figure 18: Physical space visualizations of vorticity contours (z-component) on an unseen cavity snapshot at Re=1600 (with zoom-in near the cavity region). Shown is target (P=7, top left), coarse input (P=1, middle left), SE interpolation (bottom left), a cavity-trained GNN prediction (top right), a TGV-trained GNN prediction (middle right), and a BFS-trained GNN prediction (bottom right). Blue and red indicate vorticity of 5 and -5, respectively. GNNs correspond to Model 2, 26 neighbor configurations.
Both TGV and BFS-trained models achieve impressive reconstruction capability on the cavity configuration, although there are some deviations (i.e., the contour ”streak“ just above the cavity region is not present in the target). More specifically, when looking at the vorticity contours alone, there is little qualitative difference between the cavity-trained model (the GNN baseline here) and BFS-trained model – although the BFS and cavity configurations are indeed similar, and the BFS-trained GNN has knowledge of separation dynamics and boundary layer effects, the BFS-trained model is able to recover fine-scales near the opposing cavity wall, an aspect of the flow that was unseen during training. Perhaps most interestingly, the TGV reconstructions appear to follow the target (and cavity-trained model) quite well. This is a much more challenging extrapolation task, since the TGV-trained model has effectively no knowledge of wall and separation effects. This notion is reaffirmed when observing the presence of vorticity contours at the lower wall downstream of the cavity in the TGV-trained model, a feature that is not present in either of the BFS or cavity-trained models (as well as the target).
4.4.1 Fine-Tuning on the Cavity
A natural extension to the above geometry extrapolation analysis is to investigate the degree to which fine-tuning can improve the prediction errors of BFS- and TGV-trained models on the cavity configuration.
This strategy is briefly explored here in the context of the multiscale GNN (Model 2) configuration. To investigate fine-tuning, a subset of the total parameters in either the BFS or TGV models GNN are further trained (i.e., fine-tuned) on smaller cavity dataset, while all other parameters are kept frozen. Here, the parameter subset that is fine-tuned comes from the coarse-scale processor (CSP), which contains all of the coarse-scale message passing layers, while keeping the parameters in the fine-scale processor, encoder, and decoder layers frozen. This specific fine-tuning strategy is motivated by the universality of turbulent structures: the hypothesis is similar to conventional physics-based turbulence modeling intuition, in that the large-scale changes to the flow due to a configuration change require modification of the coarse-scale message passing interactions alone, and not those modeled by message passing at the finer P=7 edge lengthscales.
The effect of fine-tuning is highlighted in Fig. 19, which showcases streamwise velocity fields in the XY plane (zoomed into the cavity). The velocity fields also reveal additional detail into the zero-shot (not fine-tuned) model performance near the cavity – it is evident that the TGV model observes increased non-physical discontinuities near the separation point and shear layer. Interestingly, fine-tuning, which is carried by training each model for an additional 50 epochs on a smaller dataset of 11480 cavity flow-field elements (roughly 11% of the dataset size used to train the reference cavity GNN model), eliminates some of these non-physical artifacts in the TGV-based model. For both TGV and BFS models, fine-tuning serves to improve predictions near the rear cavity wall (gray circles in Fig. 19, which intuitively reflects the fact that the fine-tuning improvements are primarily felt in physical regimes that are not captured in the originating configurations (TGV and BFS).
Figure 19: Visualizations of streamwise velocity component in the cavity configuration (XY mid-planes cropped near the cavity are shown) for an unseen snapshot at Re=1600, showing the effect of fine-tuning the BFS- and TGV-trained models on the cavity. All GNN models correspond to Model 2, 26-neighbor configurations. Gray circles highlight opposing cavity region, where fine-tuning improvement is observed.
@@ 5 Conclusion
The goal of this work is to enable mesh-based super-resolution of fluid flows in 3D spatial configurations. A new graph neural network architecture is introduced to this end, which is designed to take as input a graph representation of a highly coarsened unresolved flow-field, and produce as output the corresponding super-resolved flow-field on a high-resolution graph in which fine-scale features exist. The graph generation procedure coincides with an element-local interpretation of flow-fields; it is consistent with strategies used in spectral element discretizations, and shares similarities with those used in finite element and finite volume discretizations. In this framework, the GNN is designed to operate not on the full mesh-based field at once, but on localized meshes of elements (or cells) directly. To facilitate super-resolution over element-local graphs, the architecture is composed of two key stages. In the first stage, a query element alongside a set number of neighboring coarse elements are encoded into a single latent graph using coarse-scale synchronized message passing over the element neighborhood (termed the coarse-scale processor). In the second stage, fine-scale message passing is carried out on the unpooled latent graph (termed the fine-scale processor).
It is emphasized that this GNN framework embeds certain inductive biases into the model. These include the use of GNN message passing, which inherently enforces the principle that nodes in close physical proximity exert a greater influence than those farther apart (i.e., leveraging strong spatial correlations), as well as the localization of message passing to element-local regions, introducing an additional level of locality in the model structure.
@@
Demonstration studies are performed wherein the GNN architecture is applied to achieve super-resolution in the Taylor-Green Vortex (TGV) and backward-facing step (BFS) configurations at Re=1600 and Re=3200, with hexahedral mesh-based data generated using the NekRS flow solver. Within this application, the goal is to move from a P=1 element discretization (highly coarsened projected DNS) to a P=7 (DNS-level) discretization. Additional geometry extrapolation and fine-tuning tests are conducted using a separate cavity configuration, to better assess the potential of TGV- and BFS-trained model applicability across different geometries.
@@
Demonstrations of TGV super-resolution focused on performance assessments from two angles: (1) comparison of a purely coarse-scale model configuration (Model 1, with 0 message passing layers in the FSP) with a multiscale configuration (Model 2, with layers added to the FSP), and (2) the effect of the input coarse element neighborhood size on the super-resolution accuracy. Although all configurations were found to competitively achieve super-resolution in time-extrapolated scenarios (with significant improvements obtained over spectral element interpolations), model errors were found to scale roughly in proportion to increase in Reynolds number. For the TGV, at Re=1600, jumping from 0 to 6 (and 26) neighbors improved results, with Model 2 achieving generally superior reconstruction quality over Model 1 counterparts. At Re=3200, performance between Models 1 and 2 were comparable, with minimal sensitivity to the number of neighbors (i.e., at Re=3200, information contained in the surrounding element neighborhood did not play much of a role in the super-resolution process). In contrast, the BFS showed noticeable improvement due to an increaase in coarse element neighborhood size at both Reynolds numbers of 1600 and 3200.
@@
Two Reynolds-extrapolation scenarios were tested: downwards-extrapolation (application of a model trained at Re=3200 on Re=1600 data) and upwards-extrapolation (application of a model trained at Re=1600 on Re=3200 data). Although models retained their ability to generate fine-scale information at an appreciable level, upwards-extrapolation presented a greater challenge for the GNN models, with models performing better in downwards-extrapolation settings (although downwards Reynolds extrapolation resulted in overprediction of energy at the fine scales). This is consistent with physical expectations, and was deemed to be a consequence of the training procedure: the element-local flow-fields at Re=3200 naturally span a wider range of encountered turbulent structures relative to those observed at Re=1600, resulting in a highly challenging upwards-extrapolation task in the Reynolds number.
Regarding geometry extrapolations, TGV- and BFS-trained models were evaluated on unseen snapshots sourced from flow over a cavity configuration. Both models were shown to achieve good reconstruction ability on the cavity, with (as physically expected) the BFS-trained models performing better than the TGV-trained ones. Fine-tuning the coarse-scale mesage passing layers on a smaller cavity dataset showed the potential for error reduction through transfer learning strategies. Although further investigation is ultimately needed to assess full extrapolation capability and fine-tuning advantages, these demonstrations highlight (a) the applicability of the proposed GNN modeling framework in handling entirely unseen mesh-based configurations, and (b) the feasibility of multi-configuration utilization of trained models. It is emphasized that, although not explored in this work, a super-resolution GNN model can be trained on element neighborhoods sourced from many different configurations on different meshes. Such multi-configuration training is an exciting direction for developing even more robust models.
@@
This study presents many more avenues for future work that can take direct advantage of the GNN architecture introduced here. As mentioned in Sec. 4.3, a promising extension to improve Reynolds number extrapolations is to source the training data from many different Reynolds numbers. Another pathway is to extend an incremental training procedure (explored to good effect in Ref. [52]) to the GNN setting, which decomposes the one-shot approach into a sequence of smaller, but easier, super-resolution increments. The advantage of such an approach is that the same GNN model can be applied at many different resolution increments, potentially providing a more robust model. This is related to the general study of mesh-agnostic behavior in GNN architectures – on top of extending the same model to different GNN configurations, enhancing training datasets to include multiple mesh resolutions (such as in the incremental approach) is a promising route. Additionally, within the one-shot context used here, investigation of the effect of model hyperparameters on generalization capability (such as the hidden channel dimensionality and number of message passing layers) can be explored in greater detail. Lastly, avenues for establishing enhanced theoretical understanding of the connection between GNN-based message passing and the super-resolution goal should be explored.